Let {an} be a sequence defined as a1 =3, an+1 = (1/5)an ,show that {an } an converges to zero.


1
Expert's answer
2021-06-11T09:49:34-0400

"Question:-\\
Let\space {a_n}\space be\space a\space sequence\space defined\space as\space a_1\space =3,\space a_{n+1}\space =\space (\frac{1}{5})a_n\space ,\\
show\space that\space {a_n\space }\space an\space converges\space to\space zero.\\
-------------------------------------
Solution:-\\
first\space of\space all,\space we\space find\space some\space terms\space of\space sequence\\
given\\
a_1\space =3,\space a_{n+1}\space =\space (\frac{1}{5})a_n\space ,\\
we\space have\space ;\\
a_1=3...(1)\\

for\space a_2\space ,\space put\space n=1\space \space in\space \space a_{n+1};\\
a_2=(\frac{1}{5})a_1\space \\
put\space a_1\space value\space from\space (1)\\
a_2=(\frac{1}{5})(3)\space \\
a_2=(\frac{3}{5})...(2)\space \\

for\space a_3\space ,\space put\space n=2\space \space in\space \space a_{n+1};\\
a_3=(\frac{1}{5})a_2\space \\
put\space a_2\space value\space from\space (2)\\
a_2=(\frac{1}{5})(\frac{3}{5})\space \\
a_2=(\frac{3}{5^2})...(3)\space \\

for\space a_4\space ,\space put\space n=3\space \space in\space \space a_{n+1};\\
a_4=(\frac{1}{5})a_3\space \\
put\space a_2\space value\space from\space (3)\\
a_4=(\frac{1}{5})(\frac{3}{5^2})\space \\
a_4=(\frac{3}{5^3})\space \\

similarly\space \\
a_{n-1}=(\frac{3}{5^{n-2}})\space \\
for\space a_n\space ,\space put\space n=(n-1)\space \space in\space \space a_{n+1};\\
a_n=(\frac{1}{5})a_{n-1}\space \\
put\space a_{n-1}\space \\
a_n=(\frac{1}{5})(\frac{3}{5^{n-2}})\space \\
a_n=(\frac{3}{5^{n-1}})\space \\


now\space we\space show\space that\space {a_n\space }\space an\space converges\space to\space zero.\\\space 
for\space check\space to\space converges\space ,\space we\space take\space limit\space of\space a_n\\
\lim\limits_{n\space \to\space \infin}\space a_n=\lim\limits_{n\space \to\space \infin}\space (\frac{3}{5^{n-1}})\space \\
\lim\limits_{n\space \to\space \infin}\space a_n=\space (\frac{3}{5^{\infin-1}})\space \\
\lim\limits_{n\space \to\space \infin}\space a_n=\space (\frac{3}{5^{\infin}})\space \\
\lim\limits_{n\space \to\space \infin}\space a_n=\space (\frac{3}{\infin})\space \\
\lim\limits_{n\space \to\space \infin}\space a_n=\space 0\space \\
hence\\
we\space say\space that,\space {a_n\space }\space an\space converges\space to\space zero.\\\space"